# SLGaussian: Fast Language Gaussian Splatting in Sparse Views

### v1

@@ 3.2 Multi-view Language Memory Bank
Construction of Language Memory Bank. Based on the multi-view consistent segmentation obtained, we use CLIP to encode the corresponding regions of the RGB images, resulting in high-dimensional language features $L_{i1}$ and $L_{i2}$ (512 dimensions in our experiments) for the two views. If the object does not exist in a particular view, the corresponding language feature is represented by a zero vector. Directly applying these high-dimensional language features to 3DGS would lead to a memory overflow in 3DGS. [33] addresses this issue by using a per-scene autoencoder to compress the high-dimensional features, which are later decoded during querying. However, this approach is not suitable for our method, as our goal is to infer the semantic field for new scenes without retraining. A simple autoencoder cannot effectively fit the wide range of language information in the real world. Additionally, the differences in language information between the two views might cause the inferred language features to neutralize, making it impossible to recover the corresponding semantics. Therefore, this dimensionality reduction method is not applicable.
We adopt a method that uses evenly distributed vector IDs in a low-dimensional space (set to 3 dimensions in our experiments) to represent the multi-view high-dimensional language features of specific objects. Based on the total number of object categories across all views, $N$, we partition the 3D space $[0,1]^{3}$ into approximately $3]{N}\rceil^{3}$ evenly spaced vectors, from which we randomly select $N$ vectors to serve as IDs. These IDs are then matched with the corresponding object mask indices in $\{S_{i}^{1\times H\times W}\}$, converting them into semantic label maps $\{M_{i}^{3\times H\times W}\}$. The reason we do not directly use 1D object mask labels as semantic labels is that the 1D space is too low-dimensional to effectively separate the labels representing different objects in the scene. Based on the above discussion, we construct the multi-view language feature memory bank:
$\{\text{ID}:\{L_{i1},L_{i2}\}\},$ (5)
@@
3.3 Open-vocabulary Querying
Based on the previously mentioned multi-view language memory bank, we can easily map all pixels $\{v\}$ in the rendered view $I$ to the corresponding multi-view CLIP language encodings $\{L_{1}^{v},L_{2}^{v}\}_{v}$. Simultaneously, for open-vocabulary query text $qry$, we can also obtain its language encoding $\phi_{qry}=\text{CLIP}(qry)$. Then we calculate the relevance score between $\text{CLIP}(qry)$ and the multi-view language features $\{L_{1}^{v},L_{2}^{v}\}_{v}$ for each pixel $v$, using the following formula:
@@ 3.3 Open-vocabulary Querying
4 Experiments
4.1 Settings
Implementation Details. For the training process, we conducted 80,000 iterations on two NVIDIA RTX 3090 GPUs with a batch size of 1. We choose to use MVSplat [6] as the base Gaussian prediction module. The Gaussian prediction model will be frozen after pre-trained weights were loaded, then we will focus on optimizing the semantic parameters prediction branch. We used a subset of the RealEstate10K dataset [52], which originally consists of video frames extracted from YouTube videos, with corresponding camera poses obtained through COLMAP by the authors of PixelSplat [3]. Following their setup, we resized all images to 256 × 256 resolution when training on the processed RealEstate10K subset containing multi-view mask segmentations. These multi-view masks, generated via video tracking and the SAM model [8, 22], provided spatially consistent supervisory signals. We computed a mean squared error (MSE) loss between the predicted semantic feature maps and the ground truth mask identifiers during training.
For inference, we provided two-view RGB images and their corresponding camera poses, skipping further retraining on datasets such as LERF [21] and 3D-OVS [28]. Instead, we directly input two RGB images and their camera poses from these datasets, using our model to predict the semantic-encoded features. In these inference experiments, we standardized the input image resolution to 416 × 576 for LERF and 3D-OVS. Our experiments on two RTX 3090 GPUs, each utilizing around 9GB of VRAM, demonstrated that scene inference with 416 × 576 resolution images takes approximately 0.3 seconds, while single-view rendering takes around 0.0006 seconds. Additionally, model inference on a single 3090 GPU consumes roughly 6GB of VRAM.
Baseline Comparison Settings. We conducted comparative experiments using several baselines, mainly focusing on LERF and LangSplat [33] for 3D object localization and query accuracy, evaluated with IoU and Localization Accuracy metrics. The IoU metric is calculated by measuring the overlap between the queried object and its ground truth mask. Additionally, we compared our method with LERF, LangSplat, and 3D-OVS on the 3D-OVS dataset for object segmentation, using IoU and mIoU metrics. The mIoU is calculated by segmenting the entire view and averaging the IoU scores across these segments.
@@ 4.1 Settings
Figure 3: Visualization of RealEstate10K dataset scenes. The images illustrate two scenes from the test set, displayed sequentially from left to right: the two-view input RGB images, the corresponding spatially consistent SAM segmentation results, the predicted semantic feature map for a novel view, and the 3D object localization results when querying with the word ”Chair” based on the semantic feature map.
4.2 Results on RealEstate10K dataset
@@
Table 1: Quantitative comparisons of 3D semantic segmentation (IoU, %) and localization accuracy (%) on the LERF dataset.
4.3 Comparisons on LERF dataset
Figure 4: Qualitative comparisons of open-vocabulary 3D object localization on the LERF and 3D-OVS datasets. The top row displays scenes from the LERF dataset, while the bottom row shows scenes from the 3D-OVS dataset. Red points indicate the model predictions, and black dashed bounding boxes denote the annotations.
@@ 4.3 Comparisons on LERF dataset
Figure 5: Qualitative comparisons of open-vocabulary 3D object segmentation on the 3D-OVS dataset.
4.4 Comparisons on 3D-OVS dataset
@@
Table 2: Quantitative comparisons of 3D semantic segmentation on the 3D-OVS dataset. IoU and mIoU scores (%) are reported for each method. The metrics for additional scenes can be found in the supplementary material.
@@
Table 3: The ablation study in teatime and covered desk scenes comparing 3D object query Chosen IoU (%) across different settings. The configurations tested include whether the Gaussian feed-forward inference model was used, indicated as Feed-forward Model; whether mask association was performed (MA); and whether the Multi-View Language Memory Bank was performed (MV-LMB).
@@
Table 4: Comparison of SLGaussian and 2D query method for novel viewpoint object query on the LERF dataset (416 x 576 resolution).
4.5 Ablation Study
Ablation study in teatime and covered desk scenes. Table 3 presents the results of our ablation study in teatime and covered desk scenes. Initially, without the feed-forward Model, Mask Association (MA), or Multi-View Language Memory Bank (MV-LMB), our model operated similarly to LangSplat and achieved poor Chosen IoU. However, when we introduced the feed-forward Model, the Chosen IoU dropped slightly due to its high consistency demands on multi-view inputs. After incorporating the MA module, the SAM segmentation achieved multi-view consistency, leading to an improvement in the Chosen IoU metric. However, consistent masks across views did not guarantee consistent language information, as the CLIP model [34] produces different encodings for the same object when viewed from different angles. It is important to note that when MV-LMB module was not used, the language features to learn were generated by the method in LangSplat. Finally, incorporating the MV-LMB module, which assigns consistent Label IDs to masks of the same object across views and binds these IDs to multi-view CLIP features, effectively enforced feature consistency. This addition significantly improved Chosen IoU, surpassing initial LangSplat results. Notably, when the MV-LMB module was not used, we relied on the same autoencoder approach as LangSplat for processing the high-dimensional CLIP encoding features. More ablation experimental results can be seen in our supplementary material.

### v3

@@ 3.2. Multi-view Language Memory Bank
Construction of Language Memory Bank. Based on the multi-view consistent segmentation obtained, we use CLIP to encode the corresponding regions of the RGB images, resulting in high-dimensional language features $L_{i1}$ and $L_{i2}$ (512 dimensions in our experiments) for the two views. If the object does not exist in a particular view, the corresponding language feature is represented by a zero vector. Directly applying these high-dimensional language features to 3DGS would lead to a memory overflow in 3DGS. (Qin et al., 2024) addresses this issue by using a per-scene autoencoder to compress the high-dimensional features, which are later decoded during querying. However, this approach is not suitable for our method, as our goal is to infer the semantic field for new scenes without retraining. A simple autoencoder cannot effectively fit the wide range of language information in the real world. Additionally, the differences in language information between the two views might cause the inferred language features to neutralize, making it impossible to recover the corresponding semantics. Therefore, this dimensionality reduction method is not applicable.
We adopt a method that uses evenly distributed vector IDs in a low-dimensional space (set to 3 dimensions in our experiments) to represent the multi-view high-dimensional language features of specific objects. Based on the total number of object categories across all views, $N$, we partition the 3D space $0,1]^{3}$ into approximately $\lceil\sqrt[3]{N}\rceil^{3}$ evenly spaced vectors, from which we randomly select $N$ vectors to serve as IDs. These IDs are then matched with the corresponding object mask indices in $\{S_{i}^{1\times H\times W}\}$, converting them into semantic label maps $\{M_{i}^{3\times H\times W}\}$. The reason we do not directly use 1D object mask labels as semantic labels is that the 1D space is too low-dimensional to effectively separate the labels representing different objects in the scene. Based on the above discussion, we construct the multi-view language feature memory bank:
(5) $\{\text{ID}:\{L_{i1},L_{i2}\}\},$
@@
3.3. Open-vocabulary Querying
Based on the previously mentioned multi-view language memory bank, we can easily map all pixels $\{v\}$ in the rendered view $I$ to the corresponding multi-view CLIP language encodings $\{L_{1}^{v},L_{2}^{v}\}_{v}$. Simultaneously, for open-vocabulary query text qryqry, we can also obtain its language encoding $\phi_{qry}=\text{CLIP}(qry)$. Then we calculate the relevance score between $\text{CLIP}(qry)$ and the multi-view language features $\{L_{1}^{v},L_{2}^{v}\}_{v}$ for each pixel $v$, using the following formula:
@@ 3.3. Open-vocabulary Querying
4. Experiments
4.1. Settings
Implementation Details. For the training process, we conducted 80,000 iterations on two NVIDIA RTX 3090 GPUs with a batch size of 1. We choose to use MVSplat (Chen et al., 2024a) as the base Gaussian prediction module. The Gaussian prediction model will be frozen after pre-trained weights were loaded, then we will focus on optimizing the semantic parameters prediction branch. We used a subset of the RealEstate10K dataset (Zhou et al., 2018), which originally consists of video frames extracted from YouTube videos, with corresponding camera poses obtained through COLMAP by the authors of PixelSplat (Charatan et al., 2024). Following their setup, we resized all images to 256 × 256 resolution when training on the processed RealEstate10K subset containing multi-view mask segmentations. These multi-view masks, generated via video tracking and the SAM model (Cheng et al., 2023; Kirillov et al., 2023), provided spatially consistent supervisory signals. We computed a mean squared error (MSE) loss between the predicted semantic feature maps and the ground truth mask identifiers during training.
For inference, we provided two-view RGB images and their corresponding camera poses, skipping further retraining on datasets such as LERF (Kerr et al., 2023) and 3D-OVS (Liu et al., 2023). Instead, we directly input two RGB images and their camera poses from these datasets, using our model to predict the semantic-encoded features. In these inference experiments, we standardized the input image resolution to 416 × 576 for LERF and 3D-OVS. Our experiments on two RTX 3090 GPUs, each utilizing around 9GB of VRAM, demonstrated that scene inference with 416 × 576 resolution images takes approximately 0.3 seconds, while single-view rendering takes around 0.0006 seconds. Additionally, model inference on a single 3090 GPU consumes roughly 6GB of VRAM.
Baseline Comparison Settings. We conducted comparative experiments using several baselines, mainly focusing on LERF and LangSplat (Qin et al., 2024) for 3D object localization and query accuracy, evaluated with IoU and Localization Accuracy metrics. The IoU metric is calculated by measuring the overlap between the queried object and its ground truth mask. Additionally, we compared our method with LERF, LangSplat, and 3D-OVS on the 3D-OVS dataset for object segmentation, using IoU and mIoU metrics. The mIoU is calculated by segmenting the entire view and averaging the IoU scores across these segments.
@@ 4.1. Settings
Figure 3. Visualization of RealEstate10K dataset scenes. The images illustrate two scenes from the test set, displayed sequentially from left to right: the two-view input RGB images, the corresponding spatially consistent SAM segmentation results, the predicted semantic feature map for a novel view, and the 3D object localization results when querying with the word ”Chair” based on the semantic feature map.
4.2. Results on RealEstate10K dataset
@@
Table 1. Quantitative comparisons of 3D semantic segmentation (IoU, %) and localization accuracy (%) on the LERF dataset.
@@
4.3. Comparisons on LERF dataset
Figure 4. Qualitative comparisons of open-vocabulary 3D object localization on the LERF and 3D-OVS datasets. The top row displays scenes from the LERF dataset, while the bottom row shows scenes from the 3D-OVS dataset. Red points indicate the model predictions, and black dashed bounding boxes denote the annotations.
@@ 4.3. Comparisons on LERF dataset
Figure 5. Qualitative comparisons of open-vocabulary 3D object segmentation on the 3D-OVS dataset.
4.4. Comparisons on 3D-OVS dataset
@@
Table 2. Quantitative comparisons of 3D semantic segmentation on the 3D-OVS dataset. IoU and mIoU scores (%) are reported for each method. The metrics for additional scenes can be found in the supplementary material.
@@
Table 3. The ablation study in teatime and covered desk scenes comparing 3D object query Chosen IoU (%) across different settings. The configurations tested include whether the Gaussian feed-forward inference model was used, indicated as Feed-forward Model; whether mask association was performed (MA); and whether the Multi-View Language Memory Bank was performed (MV-LMB).
@@
Table 4. Comparison of SLGaussian and 2D query method for novel viewpoint object query on the LERF dataset (416 x 576 resolution).
@@
4.5. Ablation Study
Ablation study in teatime and covered desk scenes. Table 3 presents the results of our ablation study in teatime and covered desk scenes. Initially, without the feed-forward Model, Mask Association (MA), or Multi-View Language Memory Bank (MV-LMB), our model operated similarly to LangSplat and achieved poor Chosen IoU. However, when we introduced the feed-forward Model, the Chosen IoU dropped slightly due to its high consistency demands on multi-view inputs. After incorporating the MA module, the SAM segmentation achieved multi-view consistency, leading to an improvement in the Chosen IoU metric. However, consistent masks across views did not guarantee consistent language information, as the CLIP model (Radford et al., 2021) produces different encodings for the same object when viewed from different angles. It is important to note that when MV-LMB module was not used, the language features to learn were generated by the method in LangSplat. Finally, incorporating the MV-LMB module, which assigns consistent Label IDs to masks of the same object across views and binds these IDs to multi-view CLIP features, effectively enforced feature consistency. This addition significantly improved Chosen IoU, surpassing initial LangSplat results. Notably, when the MV-LMB module was not used, we relied on the same autoencoder approach as LangSplat for processing the high-dimensional CLIP encoding features. More ablation experimental results can be seen in our supplementary material.
@@ 5. Conclusion
Acknowledgments
This research was funded through National Key Research and Development Program of China (Project No. 2022YFB36066), in part by the Shenzhen Science and Technology Project under Grant (KJZD20240903103210014, JCYJ20220818101001004).